The binding energy of electrons in a metal is 201 kJ/mol .Find the threshold frequency of the metal.

Please explain how & the formulas! Thanks :)
Answer #1

The energy required to eject electron from the metal is 201 kJ/ mole which is the binding energy of the metal.

To eject one electron from the metal, required energy is : 201 kJ/ Avogadro's number = 201 kJ/ (6.023 * 1023 ) = 3.34 * 10-22 kJ

3.34 * 10-24 kJ = 3.34 * 10-22 kJ * (1000 J/ 1 kJ) = 3.34 * 10-19 J

This is the threshold energy to eject one electron from the metal.

Threshold energy = Planck's constant * Threshold Frequency

So, 3.34 * 10-19 J = 6.625 * 10-34 J.s * Threshold Frequency

Or, Threshold Frequency = ( 3.34 * 10-19 / 6.625 * 10-34 ) s-1 = 5.04 * 1014 s-1